### v2

# Exact bounds for the sum of the inverse-power of element orders in non-cyclic finite groups

Abstract
Given a finite group $G$ of order $n.$ Denote the sum of the inverse-power of element orders in $G$ by $m(G).$ Let $\mathbb{Z}_{n}$ be the cyclic group of order $n.$ Suppose $G$ is a non-cyclic group of order $n$ then we show that $m(G)\geq\frac{5}{4}m(\mathbb{Z}_{n}).$ Our result improves the inequality $m(G)>m(\mathbb{Z}_{n})$ obtained by Baniasad Azad, M., and Khorsravi B. Moreover, this bound is best as for $n=4l,$ $l$ odd, there exists a group $G$ of order $n$ satisfying $m(G)=\frac{5}{4}m(\mathbb{Z}_{n})$. Moreover, we will establish that $\frac{1}{q-1}m(G)<m(\mathbb{Z}_{n}),$ where $G$ is a non-cyclic group of an odd order $n,$ and $q$ is the smallest prime dividing $n.$
1 Introduction
The structure of a group has been studied by looking at the element orders using different invariant. For example, several authors studied several structural properties of a finite group using the function $\psi(G)=\sum_{a\in G}o(a),$ where $o(a)$ denotes the order of $a\in G,$ for example see in [HMLM]. In 2017, Garonzi, Martino, and Massimiliano Patassini introduced the function $m(G)=\sum_{a\in G}\frac{1}{o(a)},$ in [GMP]. Recently, in 2023, Morteza Baniasad Azad, Behrooz Khosravi and Hamideh Rashidi studied the group theoretical properties such as solvability, supersolvability and nilpotency of a finite group $G,$ using the function $m(G),$ (cf. [BKR, Theorem $2.7,$ $2.8,$ and $2.9$ ])
In the following theorem Garonzi, Martino, and Massimiliano Patassini provided an inequality to investigate the cyclic property of a finite group $G.$
Theorem 1.1 ( cf. [GMP, Theorem $5$]).
For any finite group $G$ of order $n$, $m(G)\geq m(\mathbb{Z}_{n})$ and equality holds if and only if $G\simeq\mathbb{Z}_{n},$ where $\mathbb{Z}_{n}$ is the cyclic group of order $n.$
In this paper, we continue our study of structural properties of a finite group using $m(G).$ The following theorem is our main result, which gives an exact bound and strengthens the bound in Theorem 1.1.
Theorem 1.2.
Let $G$ be a non-cyclic group of order $n,$ and $\mathbb{Z}_{n}$ is the cyclic group of order $n.$ Then
$m(G)\geq\frac{5}{4}m(\mathbb{Z}_{n}).$
In the following proposition it is shown that there exist a finite group of order $n$ such that $m(G)=\frac{5}{4}m(\mathbb{Z}_{n}).$ Hence this lower bound is the best possible.
Proposition 1.3.
Let $n=4l,$ where $l$ is an odd integer. Then
$m(\mathbb{Z}_{n})=2m(\mathbb{Z}_{l})\hskip 22.76219ptm(\mathbb{Z}_{2l}\times\mathbb{Z}_{2})=\frac{5}{2}m(\mathbb{Z}_{l})$
and hence
$m(\mathbb{Z}_{2l}\times\mathbb{Z}_{2})=\frac{5}{4}m(\mathbb{Z}_{n}).$
In the view of Theorem 1.2 and Proposition 1.3, it is shown that, for each $n=4l,$ where $l$ is an odd, there exist a group of order $n$ satisfying $m(G)=\frac{5}{4}m(\mathbb{Z}_{n}).$
In the next theorem we provides an improvement for non-cyclic finite groups of odd order.
Theorem 1.4.
Let $G$ be a non-cyclic finite group of an odd order $n,$ and $q$ be the smallest prime dividing $n.$ Then
$m(\mathbb{Z}_{n})>\frac{1}{(q-1)}m(G).$
2 Preliminaries
In this section we will describe some instrumental results which are already developed for the function $m(G).$
Lemma 2.1 (cf. [BKR, Lemma 2.2]).
If $A\leq G$ then $m(A)\leq m(G)$ and equality holds if and only if $A=G.$
Lemma 2.2 (cf. [BKR, Lemma 2.3]).
If $A\trianglelefteq G$ then $m(G/A)\leq m(G)$ and equality holds if and only if $A=1.$
Lemma 2.3 (cf. [BKR, Lemma 2.6]).
For any two finite groups $A_{1}$ and $A_{2},$ we have $m(A_{1}\times A_{2})\geq m(A_{1})\cdot m(A_{2}).$ Moreover, equality holds if and only if $\gcd(|A_{1}|,|A_{2}|)=1.$
Lemma 2.4 (cf. [I, Theorem 2.20]).
[…]
3 Preparatory Bounds
In the following Lemma we establish some bounds for $m(\mathbb{Z}_{n}).$
Lemma 3.1.
Let $2\leq q=p_{1}<p_{2}<\cdots<p_{t}=p$ be the prime divisors of $n$ and the corresponding Sylow subgroups of $\mathbb{Z}_{n}$ are $P_{1},P_{2},\cdots P_{t}.$ Let $exp(P_{i})$ are $r_{i}$ and $r_{1}>1$. Then
$m(\mathbb{Z}_{n})<\sqrt{n}.$
Proof.
Let $n=\prod_{i=1}^{t}(p_{i}^{r_{i}}).$ Lemma 2.3 stipulates $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(m(P_{i})),$ and for each $i,$ we have $m(P_{i})=(\frac{p_{i}-1}{p_{i}}r_{i}+1).$
So, $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}}r_{i}+1)<\prod_{i=1}^{t}(\frac{1}{2}\cdot r_{i}+1).$ For each natural number $r>1,$ we have $\frac{r}{2}+1\leq 2^{r/2}.$ Using this fact, we obtain that $m(\mathbb{Z}_{n})<\prod_{i=1}^{t}2^{r_{i}/2}.$
Additionally take note of that, $\prod_{i=1}^{t}2^{r_{i}/2}\leq\prod_{i=1}^{t}(p_{i})^{r_{i}/2},$ for $p_{i}\geq 2.$
Integrating the previously mentioned observations, we get that,
$\prod_{i=1}^{t}2^{r_{i}}<\prod_{i=1}^{t}(p_{i})^{r_{i}/2}<\sqrt{n}.$
Hence, we obtain that $m(\mathbb{Z}_{n})<\sqrt{n},$ when $q$ is the smallest prime dividing n and $q\geq 2.$
In the next Lemma we try to obtain a sharper upper bound for $m(\mathbb{Z}_{n})$ considering the smallest prime dividing $n$ is $q>5.$
Lemma 3.2.
Let $5<q=p_{1}<p_{2}<\cdots<p_{t}=p$ be the prime divisors of $n$ and the corresponding Sylow subgroups of $\mathbb{Z}_{n}$ are $P_{1},P_{2},\cdots P_{t}.$ Then
$m(\mathbb{Z}_{n})<\frac{4}{5}\sqrt{n}.$
Proof.
Let $n=\prod_{i=1}^{t}(p_{i}^{r_{i}}).$ Lemma 2.3 stipulates $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(m(P_{i})),$ and for each $i,$ we have $m(P_{i})=(\frac{p_{i}-1}{p_{i}}r_{i}+1).$
So, $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}}r_{i}+1)<\prod_{i=1}^{t}(r_{i}+1).$ For each natural number $r,$ we have $r+1\leq 2^{r}.$ Using this fact, we obtain that $m(\mathbb{Z}_{n})<\prod_{i=1}^{t}2^{r_{i}}=\prod_{i=1}^{t}4^{r_{i}/2}.$
Additionally take note of that, $\prod_{i=1}^{t}4^{r_{i}/2}\leq\prod_{i=1}^{t}(\frac{4}{5}p_{i})^{r_{i}/2},$ for $p_{i}>5.$
Integrating the previously mentioned observations, we get that,
$\prod_{i=1}^{t}2^{r_{i}}<\prod_{i=1}^{t}(\frac{4}{5})^{r_{i}/2}\prod_{i=1}^{t}(p_{i})^{r_{i}/2}<\frac{4}{5}\sqrt{n}.$
Hence, we obtain that $m(\mathbb{Z}_{n})<\frac{4}{5}\sqrt{n},$ when $q$ is the smallest prime dividing n and $q>5.$
Lemma 3.3.
Let $q=p_{1}<p_{2}<\cdots<p_{t}=p$ be the prime divisors of $n$ and the corresponding Sylow subgroups of $\mathbb{Z}_{n}$ are $P_{1},P_{2},\cdots P_{t}.$ Then
$\frac{1}{p}<m(\mathbb{Z}_{n}),\hskip 22.76219pt\hbox{and}\hskip 22.76219ptm(\mathbb{Z}_{n})>\frac{\phi(n)}{n}.$
Proof.
Let $n=\prod_{i=1}^{t}(p_{i}^{r_{i}}).$ Lemma 2.3 gives $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(m(P_{i})),$ and for each $i,$ we have $m(P_{i})=(\frac{p_{i}-1}{p_{i}}r_{i}+1).$ Now for each $i,$ it is clear that $\frac{p_{i}-1}{p_{i}}r_{i}+1>\frac{p_{i}-1}{p_{i}}$ and $p_{i+1}\geq p_{i}+1.$ Hence it follows that $m(\mathbb{Z}_{n})>\frac{1}{p}.$
For the another inequality note that, $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}}r_{i}+1).$ The value of Euler’s $\phi$-function is $\phi(n)=\prod_{i=1}^{t}\phi(p_{i}^{r_{i}})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}})p_{i}^{r_{i}}.$ So, we use the fact $\frac{p_{i}-1}{p_{i}}<(\frac{p_{i}-1}{p_{i}}\cdot r_{i}+1),$ for each $i,$ and it follows that $\frac{\phi(n)}{n}<m(\mathbb{Z}_{n}).$
To obtain another lower bound for $m(\mathbb{Z}_{n}),$ we will obtain a lower bound for the Euler’s function $\phi(n)$ in the following lemma:
Lemma 3.4.
Let $p$ be the largest prime divisor and $q$ be the smallest prime divisor of an integer $n>1$ respectively. Then $\frac{\phi(n)}{n}\geq\frac{q-1}{p}.$
Proof.
Let $t$ be the number of distinct prime divisors of $n.$ If $t=1,$ we have $n=p^{r},$ for some $r\geq 1,$ and $\phi(n)=(p-1)p^{r-1}=(p-1)(n/p).$ Next assume that $t>1,$ and the proof will be done by induction on $t.$
Now $n=p^{s}\cdot m,$ where, $(p,m)=1.$ Then $m$ has exactly $t-1$ prime divisors. Hence,
$\phi(n)=\phi(p^{s})\cdot\phi(m)\geq(p-1)\frac{p^{s}}{p}\cdot\frac{m(q-1)}{p}=\frac{(p-1)(q-1)n}{pq}\geq\frac{n(q-1)}{p}.$
The last inequality holds since, $p-1\geq q.$
Lemma 3.5.
Let $p$ be the largest prime divisor and $q$ be the smallest prime divisor of an integer $n>1$ respectively. Then $m(\mathbb{Z}_{n})\geq\frac{(p-1)(q-1)}{pq}.$
Proof.
Combining the results of Lemma 3.3 and Lemma 3.4, the result follows. ∎
4 Proofs of the main results
We will first proof following two Lemmas and then we will move for Theorem 1.2. Later we will prove Theorem 1.4, which is specially for the finite groups of an odd order.
Lemma 4.1.
Let $G$ be a non-cyclic group of order $n,$ where $q$ is the smallest prime dividing $n,$ with $q>5.$ Then
$m(G)\geq\frac{5}{4}m(\mathbb{Z}_{n}).$
Proof.
Equivalent statement of theorem is: If $m(G)<\frac{5}{4}m(\mathbb{Z}_{n})$ then $G$ is a cyclic group. We will proof it by induction on the cardinality of the finite group $G.$
If $|G|=1,$ then the result is trivial. So move to the case when $|G|>1.$ Suppose that $G$ contains a non-trivial normal cyclic subgroup $N$ such that $(|N|,|G/N|)=1.$ Then $m(G/N)<m(G).$ Now,
$m(N)\cdot m(G/N)=m(N\times G/N)=m(G)<\frac{5}{4}m(\mathbb{Z}_{n})=\frac{5}{4}m(\mathbb{Z}_{|N|})\cdot m(\mathbb{Z}_{|G/N|}).$
Since, $N$ is a cyclic group, we obtain that $m(G/N)<\frac{5}{4}m(\mathbb{Z}_{|G/N|}).$ This implies $G/N$ is a cyclic subgroup of $G.$ Hence, $G$ is a cyclic group.
So we may assume that $G$ has no non-trivial normal cyclic subgroup with $(|N|,|G/N|)=1.$ Now using Lemma 3.2, we have that
$m(G)<\frac{5}{4}\cdot m(\mathbb{Z}_{n})<\sqrt{n}.$
So, there exists an element $x\in G$ such that $o(x)>\sqrt{n},$ and $|\langle x\rangle|>|G:\langle x\rangle|.$
[…]
Suppose, $(|N|,|G/N|)\neq 1.$ Hence, $G$ is an abelian group but not a cyclic group. So, $G$ has no element of of order $n.$ Let $\alpha$ be the maximum possible order of a non-identity element in $G.$ Let $b$ be the largest possible order of a non-identity, non-generating element in $\mathbb{Z}_{n}.$ Now, $m(G)<\frac{5}{4}m(\mathbb{Z}_{n}).$ So we obtain that,
$\frac{\phi(n)}{n}+\frac{n-\phi(n)}{b}<m(G)<\frac{5}{4}m(\mathbb{Z}_{n})<\frac{\phi(n)}{\alpha}\cdot\frac{5}{4}+(n-\phi(n))\cdot\frac{5}{4}.$
It implies that,
$\frac{b}{5b-4}<\frac{(n-\phi(n))}{\phi(n)}\cdot\frac{\alpha n}{(4\alpha-5n)}<0,$
as $\alpha<n.$ Then either $b<0$ or $b<\frac{4}{5}.$ But both cases are absurd, and so it is a contradiction. Hence, $(|N|,|G/N|)=1.$ This completes the proof of Lemma 4.1. ∎
Lemma 4.2.
Let $G$ be a non-cyclic group of order $n,$ where $q$ is the smallest prime dividing $n,$ with $q\geq 2,$ and $p$ is the largest prime dividing $n$ with $p\in\{2,3,5\}.$ Then
$m(G)\geq\frac{5}{4}m(\mathbb{Z}_{n}).$
Proof.
In the case, $p\leq 5,$ and $q\geq 2,$ Lemma 3.5 and Theorem 1.1 gives that, $\frac{\phi(n)}{n}\geq\frac{1}{5}.$
In a non-cyclic group $G,$ order of each non-identity elements are at most $n/2.$ So, we obtain that,
$\frac{4}{5}\cdot m(G)\geq\frac{4}{5}+\frac{4(n-1)}{5}\cdot\frac{2}{n}.$
In the cyclic group $\mathbb{Z}_{n},$ there are $\phi(n)$ number of elements of order $n,$ and the order of rest of all non-identity elements are at least $2$ Hence, we have
$m(\mathbb{Z}_{n})\leq 1+\frac{\phi(n)}{n}+\frac{(n-\phi(n)-1)}{2}=\frac{n+1}{2}-\phi(n)\cdot\frac{(n-2)}{2n}.$
It is enough to show that
$\frac{n+1}{2}-\phi(n)\cdot\frac{(n-2)}{2n}\leq\frac{4}{5}+\frac{4(n-1)}{5}\cdot\frac{2}{n}.$
So, our claim is:
$\Leftarrow\frac{4}{5}+\frac{4(n-1)}{5}\cdot\frac{2}{n}-\frac{n+1}{2}+\phi(n)\cdot\frac{(n-2)}{2n}\geq 0$
$\Leftarrow\frac{4}{5}+\frac{4(n-1)}{5}\cdot\frac{2}{n}+[\phi(n)\cdot\frac{(n-2)}{2n}-\frac{n+1}{2}]\geq 0$
$\Leftarrow\frac{4}{5}+\frac{8(n-1)}{5n}+[5\cdot\frac{(n-2)}{2}-\frac{n+1}{2}]\geq 0,$ (4.1)
and this holds for $n>\frac{11}{4},$ which is obvious.
In the case, $p\leq 3$ and $q\geq 2,$ we have, $\phi(n)\geq 3n.$ Using the similar method as in equation 4.1, one can calculate that $m(\mathbb{Z}_{n})\leq\frac{4}{5}m(G)$ holds if $n>\frac{7}{2},$ which is obvious.
In the case, $p=q=2,$ we have, $\phi(n)\geq 2n.$ Following the methods for equation 4.1 one can observe that $m(\mathbb{Z}_{n})\leq\frac{4}{5}m(G)$ holds if $n>5.$ Here notice that, $\frac{4}{5}\cdot m(\mathbb{Z}_{2}\times\mathbb{Z}_{2})=\frac{4}{5}\cdot\frac{5}{2}=m(\mathbb{Z}_{4}).$ This completes the proof of the Lemma 4.2. ∎
proof of Theorem 1.2.
Let $n=\prod_{i=1}^{t}p_{i}^{r_{i}},$ and $p_{i}>5$ for $i>3.$ Denote, $\prod_{i=1}^{3}p_{i}^{r_{i}}$ by $\alpha$ and $\prod_{i=4}^{t}p_{i}^{r_{i}}$ by $\beta.$ If $\beta=1$ then it will follow from Lemma 4.2. So, consider $\beta>1.$ Also, consider the exponent of 2 in $\alpha$ is strictly greater than $1,$ or $n$ is an odd integer. Clearly, $\gcd(\alpha,\beta)=1.$
Equivalent statement of theorem is: If $m(G)<\frac{5}{4}m(\mathbb{Z}_{n})$ then $G$ is a cyclic group. We will proof it by induction on the cardinality of the finite group $G.$
If $|G|=1,$ then the result is trivial. So move to the case when $|G|>1.$ Suppose that $G$ contains a non-trivial normal cyclic subgroup $N$ such that $(|N|,|G/N|)=1.$ Then $m(G/N)<m(G).$ Now,
$m(N)\cdot m(G/N)=m(N\times G/N)=m(G)<\frac{5}{4}m(\mathbb{Z}_{n})=\frac{5}{4}m(\mathbb{Z}_{|N|})\cdot m(\mathbb{Z}_{|G/N|}).$
Since, $N$ is a cyclic group, we obtain that $m(G/N)<\frac{5}{4}m(\mathbb{Z}_{|G/N|}).$ This implies $G/N$ is a cyclic subgroup of $G.$ Hence, $G$ is a cyclic group.
So we may assume that $G$ has no non-trivial normal cyclic subgroup with $(|N|,|G/N|)=1.$ Now using Lemma 3.1, and Lemma 3.2, we have that
$m(G)<\frac{5}{4}\cdot m(\mathbb{Z}_{n})=\frac{5}{4}\cdot m(\mathbb{Z}_{\alpha})m(\mathbb{Z}_{\beta})<\frac{5}{4}\sqrt{\alpha}\frac{4}{5}\sqrt{\beta}<\sqrt{n}.$
So, there exists an element $x\in G$ such that $o(x)>\sqrt{n},$ and $|\langle x\rangle|>|G:\langle x\rangle|.$
[…]
Now the only remaining case is $n=2\cdot r,$ where $r$ is an odd then $G$ has a non-trivial normal subgroup $N$ of order $r.$ Then:
$m(G)=m(\mathbb{Z}_{2})\cdot m(N)\geq m(\mathbb{Z}_{2})\cdot\frac{5}{4}m(\mathbb{Z}_{r}),$
where, the last inequality holds from the previous part of the proof. This completes the proof the main result. ∎
We continue with the proof of Theorem 1.4:
proof of Theorem 1.4.
In the cyclic group $\mathbb{Z}_{n},$ there are $\phi(n)$ number of elements of order $n,$ and the order of rest of all non-identity elements are atmost $\frac{n}{q}.$ Hence, we have
$m(\mathbb{Z}_{n})\geq 1+\frac{\phi(n)}{n}+(n-\phi(n)-1)\cdot\frac{q}{n}=1-\frac{\phi(n)}{n}\cdot(q-1)+\frac{(n-1)q}{n}.$
On the other hand, in a non-cyclic group $G,$ order of each non-identity elements are at least $q.$ So, we obtain that,
$\frac{m(G)}{(q-1)}\leq\frac{(n-1)}{q(q-1)}+\frac{1}{(q-1)}.$
Notice that,
$\frac{1}{q-1}+\frac{(n-1)}{q(q-1)}+\frac{\phi(n)}{n}\cdot(q-1)<\frac{1}{q-1}+\frac{(n-1)}{q(q-1)}+(q-1)<1+\frac{(n-1)q}{n},$
and the last inequality holds for $q^{2}(q-1)<n.$
For $n=q^{2}$, clearly one have
$\frac{1}{(q-1)}\cdot m(\mathbb{Z}_{q}\times\mathbb{Z}_{q})=\frac{1}{q-1}+\frac{(q+1)}{q}\leq m(\mathbb{Z}_{q^{2}}),$
as $q\geq 3.$ This completes the proof of the result. ∎
Corollary 4.3.
Let $G$ be a non-cyclic finite group of an odd order $n,$ where $q$ is the smallest prime dividing $n,$ and $\mathbb{Z}_{n}$ be the cyclic group of order $n.$ Then
$\frac{1}{(q-1)}<\frac{m(\mathbb{Z}_{n})}{m(G)}\leq\frac{4}{5}.$
Proof.
This corollary follows from Theorem 1.2 and Theorem 1.4. ∎
Acknowledgments
The author is supported by a post-doctoral fellowship provided by National Board for Higher Mathematics, India, (File no: 0204/27/(5)/2023/R&d-II/1183), during this work. She would like to thank Statistics and Mathematics Unit, Indian Statistical Institute, Bangalore center, India for providing wonderful research atmosphere and resources.