### v3

# Exact bounds for the sum of the inverse-power of element orders in non-cyclic finite abelian groups

Abstract
Given a finite abelian group $G$ of order $n,$ and denote the sum of the inverse-power of the element orders in $G$ by $m(G).$ Let $\mathbb{Z}_{n}$ be the cyclic group of order $n.$ Suppose $G$ is a non-cyclic abelian group of order $n.$ Then we show that $m(G)\geq\frac{3+t}{2+t}m(\mathbb{Z}_{n})$, where $n=2^{t}\cdot l$, $t\geq 0$ and $\gcd(2,l)=1.$ It improves the inequality $m(G)>m(\mathbb{Z}_{n})$ obtained by Azad and et al. Moreover, the obtained bound in this article is the best, as for $n=2^{t}l,$ $t\geq 0,$ $l$ odd, there exists a non-cyclic abelian group $G$ of order $n$ satisfying $m(G)=\frac{3+t}{2+t}m(\mathbb{Z}_{n})$. Later, we will obtain a lower bound for the ratio $\frac{m(\mathbb{Z}_{n})}{m(G)}$. So, we finally establish that that $\frac{1}{p-1}m(G)<m(\mathbb{Z}_{n})\leq\frac{2+t}{3+t}m(G),$ where $p$ is the largest prime dividing $n.$ These results provide a more precise understanding of the behavior of $m(G)$ relative to the cyclic case and establish optimal bounds in the context of finite abelian groups.
1 Introduction
The structure of a group has been studied by examining at the element orders using different invariants. For example, several authors have studied various structural properties of a finite group using the function $\psi(G)=\sum_{a\in G}o(a),$ where $o(a)$ denotes the order of $a\in G,$ for example see [HMLM]. In 2017, Garonzi, Martino, and Massimiliano Patassini introduced the function $m(G)=\sum_{a\in G}\frac{1}{o(a)},$ in [GMP]. Recently, in 2023, Morteza Baniasad Azad, Behrooz Khosravi and Hamideh Rashidi studied the group theoretical properties such as solvability, supersolvability and nilpotency of a finite group $G,$ using the function $m(G),$ (cf. [BKR, Theorem $2.7,$ $2.8,$ and $2.9$ ])
In the following theorem, Garonzi, Martino, and Massimiliano Patassini provided an inequality to investigate the cyclic property of a finite group $G.$
Theorem 1.1 ( cf. [GMP, Theorem $5$]).
For any finite group $G$ of order $n$, $m(G)\geq m(\mathbb{Z}_{n})$ and equality holds if and only if $G\simeq\mathbb{Z}_{n},$ where $\mathbb{Z}_{n}$ is the cyclic group of order $n.$
In this paper, we continue our study of structural properties of a finite abelian group using $m(G).$ The following theorem is our main result, which gives an exact bound and strengthens the bound in Theorem 1.1.
Theorem 1.2.
Let $G$ be a non-cyclic abelian group of order $n,$ and $\mathbb{Z}_{n}$ is the cyclic group of order $n.$ Then
$m(G)\geq\frac{3+t}{2+t}m(\mathbb{Z}_{n}),$
where $n=2^{t}\cdot l,$ $t\geq 0$, and $\gcd(2,l)=1.$
In the following proposition, it is shown that there exists a finite group of order $n$ such that this lower bound is the best possible.
Proposition 1.3.
Let $n=2^{t}l,$ where $t\geq 0,$ and $l$ is an odd integer. Then
$m(\mathbb{Z}_{n})=\frac{2+t}{2}m(\mathbb{Z}_{l})\hskip 22.76219ptm(\mathbb{Z}_{2l}\times\mathbb{Z}_{2^{t-1}})=\frac{3+t}{2}m(\mathbb{Z}_{l})$
and hence
$m(\mathbb{Z}_{2l}\times\mathbb{Z}_{2^{t-1}})=\frac{3+t}{2+t}m(\mathbb{Z}_{n}).$
In the view of Theorem 1.2 and Proposition 1.3, it is shown that, for each $n=2^{t}\cdot l,$ where $t\geq 0,$ $l$ is an odd, there exist a group of order $n$ satisfying $m(G)=\frac{3+t}{2+t}m(\mathbb{Z}_{n}).$
In the next theorem we provides an improvement for non-cyclic finite abelian groups of an odd order.
Theorem 1.4.
Let $G$ be a non-cyclic finite abelian group of an odd order $n,$ and $p$ be the largest prime dividing $n.$ Then
$m(\mathbb{Z}_{n})>\frac{1}{(p-1)}m(G).$
2 Preliminaries
In this section we will describe some instrumental results which are already developed for the function $m(G).$
Lemma 2.1 (cf. [BKR, Lemma 2.2]).
If $A\leq G$ then $m(A)\leq m(G)$ and equality holds if and only if $A=G.$
Lemma 2.2 (cf. [BKR, Lemma 2.3]).
If $A\trianglelefteq G$ then $m(G/A)\leq m(G)$ and equality holds if and only if $A=1.$
Lemma 2.3 (cf. [BKR, Lemma 2.6]).
For any two finite groups $A_{1}$ and $A_{2},$ we have $m(A_{1}\times A_{2})\geq m(A_{1})\cdot m(A_{2}).$ Moreover, equality holds if and only if $\gcd(|A_{1}|,|A_{2}|)=1.$
Lemma 2.4 (cf. [I, Theorem 2.20]).
[…]
3 Preparatory Bounds
In the following lemmas we establish some bounds for $m(\mathbb{Z}_{n}).$
Lemma 3.1.
Let $2\leq q=p_{1}<p_{2}<\cdots<p_{s}=p$ be the prime divisors of $n$ and the corresponding Sylow subgroups of $\mathbb{Z}_{n}$ are $P_{1},P_{2},\cdots P_{s}.$ Let the highest power of each $p_{i}$ in $P_{i}$ i.e., exponent of $p_{i}$ in $(P_{i})$ are $r_{i}.$Then
$m(\mathbb{Z}_{n})<\frac{2+t}{3+t}\sqrt{n},$
where $n=2^{t}\cdot l,$ $t\geq 0$, and $\gcd(2,l)=1.$
Proof.
Let $n=\prod_{i=1}^{s}(p_{i}^{r_{i}}).$ Lemma 2.3 stipulates $m(\mathbb{Z}_{n})=\prod_{i=1}^{s}(m(P_{i})),$ and for each $i,$ we have $m(P_{i})=(\frac{p_{i}-1}{p_{i}}r_{i}+1).$
So, $m(\mathbb{Z}_{n})=\prod_{i=1}^{s}(\frac{p_{i}-1}{p_{i}}r_{i}+1)<\prod_{i=1}^{s}(r_{i}+1).$ For each natural number $r\geq 1,$ we have $r+1\leq 2^{r}=4^{r/2}<(\frac{2+t}{3+t})^{r/2}(p)^{r/2},$ for all $p\geq 7.$ We will complete the rest of the proof dividing in the following cases.
case (a): $p_{1}=2,$ $r_{1}\neq 1$ or $r_{1}=0,$ and $r_{i}\geq 0$ for all $2\leq i\leq s$
Note that, $(\frac{r}{2}+1)\leq 2^{r/2}$ for $r=0,$ and for all natural number $r\geq 2.$ Also, $(\frac{2}{3}r+1)<3^{r/2}$ for all $r\geq 1$, and $(\frac{4}{5}r+1)<5^{r/2}$ for all $r\geq 1$. Now, $m(\mathbb{Z}_{n})=(\frac{r_{1}}{2}+1)(\frac{2r_{2}}{3}+1)(\frac{4r_{3}}{5}+1)\prod_{i=4}^{s}(\frac{p_{i}-1}{p_{i}}r_{i}+1)<2^{r_{1}/2}3^{r_{2}/2}5^{r_{3}/2}\prod_{i=4}^{s}(\frac{2+t}{3+t})^{r_{i}/2}(p_{i})^{r_{i}/2}.$ Hence, $m(\mathbb{Z}_{n})<\frac{2+t}{3+t}\sqrt{n}.$
case (b): $p_{1}=2,$ $r_{1}=1$ and $r_{i}\geq 0$ for all $2\leq i\leq s$
In this case, $n=2\cdot l,$ where $l$ is an odd integer. So, $m(\mathbb{Z}_{n})=m(\mathbb{Z}_{l})\cdot m(\mathbb{Z}_{2})<\frac{2}{3}\sqrt{l}(1+\frac{1}{2})=\sqrt{l}=\frac{\sqrt{n}}{\sqrt{2}}<\frac{3}{4}\sqrt{n}.$
Hence, we obtain that $m(\mathbb{Z}_{n})<\frac{2+t}{3+t}\sqrt{n},$ where $n=2^{t}\cdot l,$ $t\geq 0$, and $\gcd(2,l)=1.$
Next, in the following lemma we calculate some lower bounds for $m(\mathbb{Z}_{n}).$
Lemma 3.2.
Let $2\leq q=p_{1}<p_{2}<\cdots<p_{t}=p$ be the prime divisors of $n$ and the corresponding Sylow subgroups of $\mathbb{Z}_{n}$ are $P_{1},P_{2},\cdots P_{t}.$ Then
$\frac{1}{p}<m(\mathbb{Z}_{n}),\hskip 22.76219pt\hbox{and}\hskip 22.76219ptm(\mathbb{Z}_{n})>\frac{\phi(n)}{n},$
where $\phi(n)$ is the Euler’s phi-function.
Proof.
Let $n=\prod_{i=1}^{t}(p_{i}^{r_{i}}).$ Lemma 2.3 gives $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(m(P_{i})),$ and for each $i,$ we have $m(P_{i})=(\frac{p_{i}-1}{p_{i}}r_{i}+1).$ Now for each $i,$ it is clear that $\frac{p_{i}-1}{p_{i}}r_{i}+1>\frac{p_{i}-1}{p_{i}}$ and $p_{i+1}\geq p_{i}+1.$ Hence it follows that $m(\mathbb{Z}_{n})>\frac{1}{p}.$
For the another inequality note that, $m(\mathbb{Z}_{n})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}}r_{i}+1).$ The value of Euler’s $\phi$-function is $\phi(n)=\prod_{i=1}^{t}\phi(p_{i}^{r_{i}})=\prod_{i=1}^{t}(\frac{p_{i}-1}{p_{i}})p_{i}^{r_{i}}.$ So, we use the fact $\frac{p_{i}-1}{p_{i}}<(\frac{p_{i}-1}{p_{i}}\cdot r_{i}+1),$ for each $i,$ and it follows that $\frac{\phi(n)}{n}<m(\mathbb{Z}_{n}).$
To obtain another lower bound for $m(\mathbb{Z}_{n}),$ we will obtain a lower bound for the Euler’s phi-function, $\phi(n),$ in the following lemma:
Lemma 3.3.
Let $p$ be the largest prime divisor and $q$ be the smallest prime divisor of an integer $n>1$ respectively. Let $2\leq q=p_{1}<p_{2}<\cdots<p_{t}=p$ be the prime divisors of $n.$ Then $\frac{\phi(n)}{n}\geq\frac{q-1}{p}.$
Proof.
Let $t$ be the number of distinct prime divisors of $n.$ If $t=1,$ we have $n=p^{r},$ for some $r\geq 1,$ and $\phi(n)=(p-1)p^{r-1}=(p-1)(n/p).$ Next assume that $t>1,$ and the proof will be done by induction on $t.$
Now $n=p^{s}\cdot m,$ where, $(p,m)=1.$ Then $m$ has exactly $t-1$ prime divisors. Hence,
$\phi(n)=\phi(p^{s})\cdot\phi(m)\geq(p-1)\frac{p^{s}}{p}\cdot\frac{m(q-1)}{p_{t-1}}=\frac{(p-1)(q-1)n}{pp_{t-1}}\geq\frac{n(q-1)}{p}.$
The last inequality holds since, $p-1\geq p_{t-1}.$
Lemma 3.4.
Let $p$ be the largest prime divisor and $q$ be the smallest prime divisor of an integer $n>1$ respectively. Then $m(\mathbb{Z}_{n})\geq\frac{(q-1)}{p}.$
Proof.
Combining the results of Lemma 3.2 and Lemma 3.3, the result follows. ∎
4 Proofs of the main results
We will first proof Theorem 1.2. Later, we will prove Theorem 1.4, which specially applies to finite groups of odd order.
Proof.
proof of Theorem 1.2
Equivalent statement of theorem is: If $m(G)<\frac{3+t}{2+t}m(\mathbb{Z}_{n})$ then $G$ is a cyclic group. We will proof it by induction on the cardinality of the finite group $G.$
If $|G|=1,$ then the result is trivial. So move to the case when $|G|>1.$ Suppose that $G$ contains a non-trivial normal cyclic subgroup $N$ such that $(|N|,|G/N|)=1.$ Then $m(G/N)<m(G).$ Now, Lemma 2.3, and Theorem 1.1 gives that,
$m(N)\cdot m(G/N)=m(N\times G/N)\leq m(G)<\frac{3+t}{2+t}m(\mathbb{Z}_{n})=\frac{3+t}{2+t}m(\mathbb{Z}_{|N|})\cdot m(\mathbb{Z}_{|G/N|}).$
Since, $N$ is a cyclic group, $N\cong\mathbb{Z}_{|N|}.$ So we obtain that $m(G/N)<\frac{3+t}{2+t}m(\mathbb{Z}_{|G/N|}).$ This implies $G/N$ is a cyclic subgroup of $G.$ Hence, $G$ is a cyclic group.
So we may assume that $G$ has no non-trivial normal cyclic subgroup with $(|N|,|G/N|)=1.$ Now using Lemma 3.1, we have that
$m(G)<\frac{3+t}{2+t}\cdot m(\mathbb{Z}_{n})<\sqrt{n}.$
So, there exists an element $x\in G$ such that $o(x)>\sqrt{n},$ and $|\langle x\rangle|>|G:\langle x\rangle|.$
[…]
Suppose $(|N|,|G/N|)\neq 1.$ Hence, $G$ is an abelian group but not a cyclic group. Thus, $G$ has no element of of order $n.$ The least possible order of a non-identity element in $G$ is $q.$ Then
$m(G)\leq\frac{n-1}{q}+1.$
In $\mathbb{Z}_{n},$ the maximum possible value of a non-identity, non-generating element is $\frac{n}{q}.$ So,
$\frac{\phi(n)}{n}+\frac{(n-\phi(n))q}{n}<1+\frac{\phi(n)}{n}+\frac{(n-\phi(n)-1)q}{n}\leq m(\mathbb{Z}_{n}).$
Theorem 1.1 and our assumption gives that,
$m(\mathbb{Z}_{n})\leq m(G)<\frac{3+t}{2+t}m(\mathbb{Z}_{n})\leq\frac{3+t}{2+t}m(G).$
So we obtain that,
$\frac{\phi(n)}{n}+\frac{(n-\phi(n))q}{n}<\frac{3+t}{2+t}\cdot\frac{n-1}{q}+\frac{3+t}{2+t}$
$\Rightarrow\frac{\phi(n)}{n}\cdot(1-q)+q<\frac{3+t}{2+t}\cdot\frac{(n-1+q)}{q}$
$\Rightarrow\frac{\phi(n)}{n}\cdot(1-q)<\frac{((3+t)n-(3+t)+(3+t)q-(2+t)q^{2})}{(2+t)q}$
$\Rightarrow\frac{\phi(n)}{n}\cdot\frac{(2+t)q}{((3+t)n-(3+t)+(3+t)q-(2+t)q^{2})}<\frac{1}{1-q}<0$
$\Rightarrow(3+t)n-(3+t)+(3+t)q-(2+t)q^{2}<0$
$\Rightarrow(3+t)n<(3+t)-(3+t)q+(2+t)q^{2}<(3+t)+(3+t)q^{2}$
$\Rightarrow n<1+q^{2}.$
So it is a contradiction except for $n=q^{2}.$ Hence, except for the case $n=q^{2}$ we obtain $(|N|,|G/N|)=1,$ and $G$ is a cyclic group of order $n.$ To complete the proof, we need to check the case where $n=q^{2}.$ Note that $m(\mathbb{Z}_{q^{2}})=1+\frac{2(q-1)}{q},$ and $m(\mathbb{Z}_{q}\times\mathbb{Z}_{q})=(1+\frac{q^{2}-1}{q}).$ As, $q\geq 2,$ it is obvious that $m(\mathbb{Z}_{q^{2}})\leq\frac{2+t}{3+t}m(\mathbb{Z}_{q}\times\mathbb{Z}_{q}).$ This completes the proof of Theorem 1.2. ∎
We continue with the proof of Theorem 1.4:
proof of Theorem 1.4.
Suppose, $\frac{m(\mathbb{Z}_{n})}{m(G)}\leq\frac{1}{p-1}$. Now, Theorem 1.2 exhibits the best upper bound for $\frac{m(\mathbb{Z}_{n})}{m(G)}$ is $\frac{2+t}{3+t}.$ Hence, $\frac{1}{p-1}\geq\frac{2+t}{3+t}$, which implies that $p\leq 2+\frac{1}{2+t}.$ This is a contradiction expect the case $n=p^{2}.$
For $n=p^{2}$, clearly one have
$\frac{1}{(p-1)}\cdot m(\mathbb{Z}_{p}\times\mathbb{Z}_{p})=\frac{1}{p-1}+\frac{(p+1)}{p}\leq m(\mathbb{Z}_{p^{2}}),$
as $p\geq 3.$ This completes the proof of the result. ∎
Corollary 4.1.
Let $G$ be a non-cyclic finite abelian group of order $n,$ where $n=2^{t}\cdot l$ with $\gcd(2,l)=1,$ and $p$ is the largest prime dividing $n.$ Let $\mathbb{Z}_{n}$ be the cyclic group of order $n.$ Then
$\frac{1}{(p-1)}<\frac{m(\mathbb{Z}_{n})}{m(G)}\leq\frac{2+t}{3+t}.$
Proof.
This corollary follows from Theorem 1.2 and Theorem 1.4. ∎
Acknowledgments
The author is supported by a post-doctoral fellowship provided by the University of Antwerp,Belgium. She would like to thank the Department of Mathematics, University of Antwerp for providing a wonderful research atmosphere and resources.